OpenStudy (anonymous):

At the end of the winter, all snowboards at a sporting goods store were on sale for 2/3 off. What was the sale price of a snowboard that originally cost $240?

OpenStudy (anonymous):

240 - 2/3 x 240 = 240 - 160 = 80